OpenStudy (anonymous):

pat and ted traveled 79times to different places, they determined that pat had traveled five less than three times the number of times ted made. how many times did each traveled?

OpenStudy (whpalmer4):

Write the equations for the information provided. Let p be the number of times Pat traveled, and t be the number of times Ted traveled. Together, they traveled 79 times: t + p = 79 Pat traveled 5 less than 3 times the number of times Ted traveled: p = 3*t - 5 Now solve those two equations by substitution or elimination (I would suggest substitution as you've already got a formula for one of them that requires no manipulation to substitute).

OpenStudy (anonymous):

jeff has one-fourth as many books as alan. together they have 60 books. how many books does each have?

OpenStudy (mathstudent55):

Let a = number of books alan has Let j = number of books jeff has "jeff has one-fourth as many books as alan" j = (1/4)a "together they have 60 books" j + a = 60 Solve the equations simultaneously for a and j.